# LEO Constellations as a Decentralized GNSS Network: Optimizing PNT Corrections in Space

### v1

@@ I Introduction
Low Earth orbit (LEO) satellites have become a transformative platform for communications, navigation, remote sensing, Earth observation, and numerous other space applications [9681631, 9110855, 9502642, reid2025xona, 10945753, 9984725]. Many of these applications require accurate knowledge of the satellite’s position and clock state, and onboard Global Navigation Satellite System (GNSS) receivers have become a widely used solution for orbit determination and time synchronization [selvan2023precise, allahvirdi2022precise, 4319208, montenbruck2025gnss]. Currently, most LEO-onboard GNSS processing still relies on standalone positioning using broadcast ephemerides embedded within GNSS signals, which provides only meter-level accuracy due to the limited precision of the broadcast products. Missions requiring higher accuracy instead depend on precise GNSS products generated by International GNSS Service (IGS) analysis centers and derived from terrestrial continuously operating reference station (CORS) networks, which provide decimeter to centimeter accuracy [selvan2023precise, montenbruck2005rapid]. However, such performance is typically achieved only in post-processed orbit determination, since most LEO spacecraft cannot maintain continuous ground connectivity and the latency of precise products limits their usability for real-time operations [selvan2023precise, allahvirdi2022precise]. Consequently, broadcast ephemerides offer real-time availability but limited positioning accuracy, whereas precise products deliver substantially higher accuracy at the cost of latency and intermittent accessibility. An increasing number of LEO missions, including location-enabled communication systems, formation flying, positioning, navigation, and timing (PNT) services, satellite altimetry, and spaceborne sensing, now require real-time precise orbit and clock determination [selvan2023precise]. Bridging this accuracy-latency gap requires a new architecture capable of delivering both high precision and real-time availability.
@@
Ground-based GNSS networks demonstrate the effectiveness of cooperative processing: observations from global CORS networks are jointly processed at centralized analysis centers to generate precise satellite clocks and orbits, atmospheric corrections, and bias models, which underpin high-precision positioning techniques such as precise point positioning (PPP) [heroux1995gps, zumberge1997precise] and network-assisted PPP real-time kinematic (PPP-RTK) [wabbena2005ppp]. Modern LEO constellations equipped with onboard GNSS receivers can, in principle, be viewed as a spaceborne analogue of such networks, in which hundreds or thousands of rapidly moving satellites collectively provide a global and geometry-rich set of GNSS observations, as conceptually illustrated in Figure 1. However, existing LEO onboard GNSS receivers typically operate independently: GNSS-derived information is not exchanged among satellites, orbit and clock states are not estimated at the constellation level, and cooperative onboard processing across the constellation remains largely unexplored [selvan2023precise]. Inspired by the success of terrestrial CORS networks, this work examines whether a LEO constellation can operate cooperatively to jointly estimate its orbit and clock states and to generate network-derived GNSS correction information directly onboard. Such a system could extend network-based GNSS processing to the space segment and enable the onboard generation of network-derived GNSS satellite clock corrections, which may serve as a supporting component for high-precision positioning applications, particularly in remote or otherwise challenging environments.
@@
LEO provides a particularly favorable environment for spaceborne GNSS-based inference. Operating above the dense layers of the atmosphere, LEO satellites are effectively free from tropospheric delay, experience substantially reduced ionospheric effects due to shorter signal propagation paths, and are largely unaffected by multipath [hofmann2008gnss, allahvirdi2022precise]. As a result, GNSS measurements collected in LEO are considerably cleaner than those obtained by ground-based receivers. In addition, the high orbital velocity of LEO spacecraft induces strong Doppler signatures and rapidly varying observation geometry, both of which enhance measurement observability and render Doppler a highly informative and complementary observable [palomo2019space]. The scale of modern LEO constellations further amplifies these advantages, as thousands of rapidly moving satellites naturally provide dense global coverage and rich geometric diversity.
@@
With sufficient connectivity, decentralized estimation offers a pathway to globally consistent solutions without forwarding raw measurements to a centralized analysis center. By shifting computation toward the data sources, such approaches reduce backhaul load, lower latency, and enable inherent parallelism with more efficient resource utilization [sayed2014diffusion, 4118472, 6994854, mcmahan2017communication]. Beyond these benefits, decentralized architectures alleviate communication and computational burdens, improve scalability, enhance resilience to node or link failures, preserve data locality, and support higher update rates as observation density increases [yang2019survey]. These properties make decentralized processing a promising direction for large, dynamic LEO constellations operating as cooperative GNSS observation networks. Nevertheless, achieving decentralized estimation that ensures system-wide coordination remains challenging in large-scale, communication-constrained LEO networks.
@@ III Estimability and Identifiability Analysis of Model Parameters
The observation equations introduced in Section II inherently lead to a rank-deficient estimation problem: the design matrix $\mathbf{A}$ is not full rank, meaning that only certain linear combinations of the parameters are identifiable, while their absolute values are not. For a networked architecture, this issue is particularly critical, as all nodes must adopt a common set of physically meaningful estimable parameters; otherwise, different nodes would converge to mutually inconsistent solutions. To address this challenge, we adopt the $\mathcal{S}$-system theory to rigorously characterize the identifiable parameter combinations of the proposed spaceborne GNSS network [baarda1973s, teunissen1985zero, odijk2016estimability].
@@ IV-A Well-defined Decentralized Formulation
However, such a centralized architecture is challenging for large LEO constellations. Transmitting full datasets to a fusion center and forming global normal equations does not scale with constellation size or observation duration, and introduces prohibitive communication, latency, and robustness challenges. These limitations motivate a decentralized formulation in which each LEO node performs local computation while exchanging only compact information with its neighbors.
@@ IV-B Momentum-Accelerated Gradient Tracking with Orbit Dynamics
We introduce a communication-efficient method to solve (13). Building upon the GT framework [nedic2017achieving], which enables decentralized optimization of coupled local objectives by tracking the network-wide gradient using only local information exchange, we incorporate Heavy-ball momentum to accelerate convergence. This acceleration reduces the communication overhead required to reach a target accuracy. As illustrated in Figure 2, we characterize the inter-satellite communication topology of LEO constellations as exhibiting distinct behaviors across two time scales: on a small time scale, the network topology remains effectively invariant, as communication links persist despite the continuous motion of the satellites; on a large time scale, the topology evolves as satellites gradually drift out of range of previous neighbors and establish new connections. Consequently, since the constellation operates iteratively through these regimes, we model the optimization scheme over a piecewise-invariant graph.
@@
Under this formulation, (15a) implements the Heavy-ball momentum update $\mathbf{v}_{l}^{k}$ with coefficient $\theta\in[0,1)$; (15b) performs a $\gamma$ descent step combined with a $\mathcal{K}$-round consensus operator $\mathcal{C}(\cdot)$ to obtain $\mathbf{z}_{l}^{k+1}$; (15c) computes the closed-form update of the local auxiliary variable $\mathbf{x}_{l}$ given $\mathbf{z}_{l}^{k+1}$ (as derived in [zheng2025decentralizedgnss]); and (15d) updates the gradient tracker $\mathbf{g}_{l}^{k+1}$. The complete iterative procedure is summarized in Algorithm 1. For simplicity, we assume that each snapshot $\mathcal{G}^{(t)}$ remains dominant for the same period length.
@@
Although multi-round consensus increases the communication cost per iteration, prior works have significantly reduced this overhead via low-bit quantization and error-feedback mechanisms [roula, zheng2025quantitative]. In such approaches, the multi-round consensus can be interpreted as applying a finite-impulse-response (FIR) graph filter implemented through repeated local exchanges [vlaski2023networked, 9496112]. Moreover, performing consensus-only steps (i.e., mixing without additional gradient evaluations) can reduce the wall-clock time by mitigating straggler effects, which is particularly beneficial when the communication window is short.
@@ V Numerical Results
Figure 3 and Figure 4 compare the orbit determination and time synchronization performance under three processing strategies: standalone per-receiver processing, network-based processing with float ambiguities, and network-based processing with integer-fixed ambiguities. As shown in Figure 3, standalone processing yields a LEO orbit determination error of $2.95$ m, reflecting the limited accuracy of individual receivers. By jointly estimating shared GNSS satellite clock states across the constellation, the network-based float solution significantly improves the orbit accuracy to $0.12$ m. When integer ambiguity resolution is enabled using the LAMBDA method [teunnissen1995least], the network-fixed solution further reduces the orbit error to $0.06$ m, achieving centimeter-level accuracy. Figure 4 shows the corresponding time synchronization performance. The standalone solution exhibits a timing error of $7.95$ ns. For the network-based solutions, the reported timing errors correspond to the network-consistent clock states. Network-based float processing achieves a timing accuracy of approximately $0.21$ ns, which is further improved to $0.11$ ns after integer ambiguity fixing. Overall, the results demonstrate that network-based onboard GNSS processing enables substantial gains in both orbit determination and time synchronization, with integer ambiguity resolution provides further performance improvements.

### v2

@@ I Introduction
Low Earth orbit (LEO) satellites have become a transformative platform for communications, navigation, remote sensing, Earth observation, and numerous other space applications [7, 29, 8, 20, 6, 9]. Many of these applications require accurate knowledge of the satellite’s position and clock state, and onboard Global Navigation Satellite System (GNSS) receivers have become a widely used solution for orbit determination and time synchronization [22, 1, 15, 12]. Currently, most LEO-onboard GNSS processing still relies on standalone positioning using broadcast ephemerides embedded within GNSS signals, which provides only meter-level accuracy due to the limited precision of the broadcast products. Missions requiring higher accuracy instead depend on precise GNSS products generated by International GNSS Service (IGS) analysis centers and derived from terrestrial continuously operating reference station (CORS) networks, which provide decimeter to centimeter accuracy [22, 11]. However, such performance is typically achieved only in post-processed orbit determination, since most LEO spacecraft cannot maintain continuous ground connectivity and the latency of precise products limits their usability for real-time operations [22, 1]. Consequently, broadcast ephemerides offer real-time availability but limited positioning accuracy, whereas precise products deliver substantially higher accuracy at the cost of latency and intermittent accessibility. An increasing number of LEO missions, including location-enabled communication systems, formation flying, positioning, navigation, and timing (PNT) services, satellite altimetry, and spaceborne sensing, now require real-time precise orbit and clock determination [22]. Bridging this accuracy-latency gap requires a new architecture capable of delivering both high precision and real-time availability.
@@
Ground-based GNSS networks demonstrate the effectiveness of cooperative processing: observations from global CORS networks are jointly processed at centralized analysis centers to generate precise satellite clocks and orbits, atmospheric corrections, and bias models, which underpin high-precision positioning techniques such as precise point positioning (PPP) [4, 33] and network-assisted PPP real-time kinematic (PPP-RTK) [26]. Modern LEO constellations equipped with onboard GNSS receivers can, in principle, be viewed as a spaceborne analogue of such networks, in which hundreds or thousands of rapidly moving satellites collectively provide a global and geometry-rich set of GNSS observations, as conceptually illustrated in Figure 1. However, existing LEO onboard GNSS receivers typically operate independently: GNSS-derived information is not exchanged among satellites, orbit and clock states are not estimated at the constellation level, and cooperative onboard processing across the constellation remains largely unexplored [22]. Inspired by the success of terrestrial CORS networks, this work examines whether a LEO constellation can operate cooperatively to jointly estimate its orbit and clock states and to generate network-derived GNSS correction information directly onboard. Such a system could extend network-based GNSS processing to the space segment and enable the onboard generation of network-derived GNSS satellite clock corrections, which may serve as a supporting component for high-precision positioning applications, particularly in remote or otherwise challenging environments.
@@
LEO provides a particularly favorable environment for spaceborne GNSS-based inference. Operating above the dense layers of the atmosphere, LEO satellites are effectively free from tropospheric delay, experience substantially reduced ionospheric effects due to shorter signal propagation paths, and are largely unaffected by multipath [5, 1]. As a result, GNSS measurements collected in LEO are considerably cleaner than those obtained by ground-based receivers. In addition, the high orbital velocity of LEO spacecraft induces strong Doppler signatures and rapidly varying observation geometry, both of which enhance measurement observability and render Doppler a highly informative and complementary observable [18]. The scale of modern LEO constellations further amplifies these advantages, as thousands of rapidly moving satellites naturally provide dense global coverage and rich geometric diversity.
@@
With sufficient connectivity, decentralized estimation offers a pathway to globally consistent solutions without forwarding raw measurements to a centralized analysis center. By shifting computation toward the data sources, such approaches reduce backhaul load, lower latency, and enable inherent parallelism with more efficient resource utilization [21, 17, 30, 10]. Beyond these benefits, decentralized architectures alleviate communication and computational burdens, improve scalability, enhance resilience to node or link failures, preserve data locality, and support higher update rates as observation density increases [28]. These properties make decentralized processing a promising direction for large, dynamic LEO constellations operating as cooperative GNSS observation networks. Recent studies have demonstrated the potential of decentralized estimation in ground-based receiver networks [32]. Nevertheless, a principled treatment of system-wide coordination under decentralized architectures remains lacking in large-scale, communication-constrained LEO networks.
@@ III Estimability and Identifiability Analysis of Model Parameters
The observation equations introduced in Section II inherently lead to a rank-deficient estimation problem: the design matrix $\mathbf{A}$ is not full rank, meaning that only certain linear combinations of the parameters are identifiable, while their absolute values are not. For a networked architecture, this issue is particularly critical, as all nodes must adopt a common set of physically meaningful estimable parameters; otherwise, different nodes would converge to mutually inconsistent solutions. To address this challenge, we adopt the $\mathcal{S}$-system theory to rigorously characterize the identifiable parameter combinations of the proposed spaceborne GNSS network [2, 24, 16].
@@ IV-A Well-defined Decentralized Formulation
However, such a centralized architecture is challenging for large-scale LEO constellations. Currently, no dedicated spaceborne computing center exists, and deploying and operating such an infrastructure would entail substantial cost and system complexity. Even if a fusion node were available, transmitting raw observations to a central processor and forming global normal equations does not scale with the constellation size or observation duration, and would incur prohibitive communication overhead, increased latency, and limited robustness. These limitations motivate a decentralized formulation, in which each LEO node performs local computation while exchanging only compact information with its neighboring satellites.
@@ IV-B Momentum-Accelerated Gradient Tracking with Orbit Dynamics
We introduce a communication-efficient method to solve (13). Building upon the GT framework [14], which enables decentralized optimization of coupled local objectives by tracking the network-wide gradient using only local information exchange, we incorporate Heavy-ball momentum to accelerate convergence. This acceleration reduces the communication overhead required to reach a target accuracy. As illustrated in Figure 2, we characterize the inter-satellite communication topology of LEO constellations as exhibiting distinct behaviors across two time scales: on a small time scale, the network topology remains effectively invariant, as communication links persist despite the continuous motion of the satellites; on a large time scale, the topology evolves as satellites gradually drift out of range of previous neighbors and establish new connections. Consequently, since the constellation operates iteratively through these regimes, we model the optimization scheme over a piecewise-invariant graph.
@@
Under this formulation, (15a) implements the Heavy-ball momentum update $\mathbf{v}_{l}^{k}$ with coefficient $\theta\in[0,1)$; (15b) performs a $\gamma$ descent step combined with a $\mathcal{K}$-round consensus operator $\mathcal{C}(\cdot)$ to obtain $\mathbf{z}_{l}^{k+1}$; (15c) computes the closed-form update of the local auxiliary variable $\mathbf{x}_{l}$ given $\mathbf{z}_{l}^{k+1}$ (as derived in [32]); and (15d) updates the gradient tracker $\mathbf{g}_{l}^{k+1}$. The complete iterative procedure is summarized in Algorithm 1. For simplicity, we assume that each snapshot $\mathcal{G}^{(t)}$ remains dominant for the same period length.
@@
Although multi-round consensus increases the communication cost per iteration, prior works have significantly reduced this overhead via low-bit quantization and error-feedback mechanisms [13, 31]. In such approaches, the multi-round consensus can be interpreted as applying a finite-impulse-response (FIR) graph filter implemented through repeated local exchanges [25, 3]. Moreover, performing consensus-only steps (i.e., mixing without additional gradient evaluations) can reduce the wall-clock time by mitigating straggler effects, which is particularly beneficial when the communication window is short.
@@ V Numerical Results
Figure 3 and Figure 4 compare the orbit determination and time synchronization performance under three processing strategies: standalone per-receiver processing, network-based processing with float ambiguities, and network-based processing with integer-fixed ambiguities. As shown in Figure 3, standalone processing yields a LEO orbit determination error of $2.95$ m, reflecting the limited accuracy of individual receivers. By jointly estimating shared GNSS satellite clock states across the constellation, the network-based float solution significantly improves the orbit accuracy to $0.12$ m. When integer ambiguity resolution is enabled using the LAMBDA method [23], the network-fixed solution further reduces the orbit error to $0.06$ m, achieving centimeter-level accuracy. Figure 4 shows the corresponding time synchronization performance. The standalone solution exhibits a timing error of $7.95$ ns. For the network-based solutions, the reported timing errors correspond to the network-consistent clock states. Network-based float processing achieves a timing accuracy of approximately $0.21$ ns, which is further improved to $0.11$ ns after integer ambiguity fixing. Overall, the results demonstrate that network-based onboard GNSS processing enables substantial gains in both orbit determination and time synchronization, with integer ambiguity resolution provides further performance improvements.